A boy pushes a 9.0kg object a distance of 7.0m along a level floor at constant velocity by pushing horizontally on it. The coefficient of kinetic friction between the object and the floor is 0.20. Calculate the work done on the object by the force the boy apply. 



1
Expert's answer
2021-06-03T13:57:59-0400

The block is moving at a constant velocity.

So the Work done by the boy will be equal to the work done by kinetic friction.

Work done by kinetic friction = (Kinetic frictional force) × ( displacement).

=

"(\mu k) \times(N)\times d""0.20 \times (9\times 9.8)\times7""123.48 Joule"

Thus work done on the object by the force the boy apply is 123.48 Joule.


* If value of g is taken as 10

Then the answer is 0.2×9×10×7=126 Joule.